General Physics

a) The density of hollow sphere of inner and outer radii of R1 and R2 is 8g/cc.If it is just immersed in water the find the ratio of the radii. b) Two liquids A and B in different test tubes of same length and negligible mass just immerse in water. The length of liquid column of A and B are L/4 and L/3 respectively. Find their density ratio. (L = length of the tube.)

Ajit Siddharth
15 Years agoGrade

3 Answers

Chetan Mandayam Nayakar
15 Years ago

let R2=kR1, 8*(4pi/3)(R1)3(k3-1)=(4pi/3)(R1)3k3, 8k3-8=k3,k=7k3=8,k=2/(7)1/3

(DA)gL/4=(DB)gL/3, DA/DB=4/3

'Win exciting gifts by answering the questions on Discussion Forum.'

AJIT AskiitiansExpert-IITD
15 Years ago

Dear Ajit ,

a) just immersed means that the weight of liquid displaced by it is just equal to its weight

       mass of sphere  =  volume * density

                                   =  4/3*∏* (R23-  R13) * 8  =  mass of water displaced   =  =  4/3*∏* (R23) * 1

             so , 7R2=8  R13    so  , R1/R2 =  (7/8)1/3

b) if the test tube are of equal lengths , then they must displace same amount of water and they both just immerse so their mass must also be equal

                     AL/4 * D1  =  AL/3*D2

                   D1/D2 =  4 /3

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
                                                                     

Win exciting gifts by answering the questions on Discussion Forum.

 

                                                                    
                                                                     
                                             
 





 

 

 

Chetan Mandayam Nayakar
15 Years ago

Dear Ajit, 

8*(4pi/3)(R23-R13)=(4pi/3)R23, 7R23=8R13,R2/R1=2/(71/3)

let dA and dB be the densities of the two liquids, and let a and b be the cross-sectional areas of A and B respectively.w=density of water

 dA(L/4)ag=Lawg, dA/4=w. similarly, we get dB/3=w, therefore dA/dB=4/3

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras